OpenStudy (anonymous):

What is the absolute value of the number? 3+i

OpenStudy (primeralph):



OpenStudy (primeralph):

you mean magnitude or norm.

OpenStudy (mathstudent55):

z = a + bi |z| = sqrt(a^2 + b^2)

OpenStudy (primeralph):